1- Simplify the following Boolean expressions to a minimum number of literals:

  1. BC+ BC'+ BA

  2. (A + C)(AD + AD) + AC + C:

  3. xyz + x’y + xyz’ 

  4. A(A + B) + (B + AA)(A + B):
Answer #1

3. xyz + x′y + xyz′ = xy(z+z′) + x′y = xy +x′y = y(x+x′)=y 

answered by: casio
Answer #2
BC+BC'+BA= B(C+C')+BA=B+BA=B(A+1)=B.1=B
source: Digital electronics
answered by: عبد الرحمن